Les and Gil drove a total of 733 miles in 11.6 hours. Les drove the first part of the trip and averaged 55 mph. Gilbert drove the remainder of the trip and averagerd 65 mph. How long did Les drive?
Answer by checkley77(12569)

55X+65(11.6-X)=733
55X+754-65X=733
-10X=733-754
-10X=-21
X=-21/-10
X=2.1 HOURS FOR LES' DRIVING TIME.
11.6-2.1=9.5 HOURS FOR gILS' DRIVING TIME.
PROOF:
.55*2.1+65*9.5=733
115.5+617.5=733
733=733